Given a spinner with the numbers one through eight, what is the probability that you will spin an even number or a number greater than four?
A) 1/4
B) 1/2
C) ¾
D) 1

Answer by stanbon(75887)

Given a spinner with the numbers one through eight, what is the probability that you will spin an even number or a number greater than four?
P(even or number>4) = P(even) + P(number>4)-P(even AND number>4)
=P(2,4,6,8) + P(5,6,7,8) - P (6,8)
= 4/8 + 4/8 - 2/8 = 6/8 = 3/4
-----------------------------------
Cheers,
Stan H.
A) 1/4
B) 1/2
C) ¾
D) 1